# Learning to summarize user information for personalized reinforcement learning from human feedback

### v2

@@ 5 Experiments
5.1 Domains
We evaluate on three benchmark tasks from  Poddar et al. [18], along with a challenging variant that tests out-of-distribution user contexts, as well as real user datasets.
Pets. In this dataset, 50% of users prefer to discuss dogs and 50% prefer cats, so a single reward model will fail to accurately capture user preferences. To test the robustness of different methods with new conversation topics and users, we create a new Pets OOD (Out-Of-Distribution) dataset, which consists of users who prefer talking about rabbits or birds, as opposed to cats and dogs. Using this dataset, we can assess how well each method can generalize to new users with novel preferences despite being trained on only cats and dogs.
UltraFeedback. Poddar et al. [18] construct two domains from the scaled-up preference dataset UltraFeedback [5]. In UltraFeedback-P2, users prefer either helpfulness or honesty, and in UltraFeedback-P4, users choose a subset of four possible attributes of an LLM assistant: helpfulness, honesty, truthfulness, and instruction-following. In both domains, personalized reward models need to infer different user types based on past conversation data. For example, some users may prioritize helpfulness and honesty, while others may value instruction-following above all other attributes. Following  Poddar et al. [18], we also focus on a controversial subset of the dataset by removing samples where all users agree on the chosen response.
@@ 5.1 Domains
Evaluation with real user data (PRISM). This real-world user dataset consists of open-ended conversations from 1,500 participants across 75 countries and 20 LLMs [12]. Each user engages in up to six conversations on topics of their choice, loosely guided by one of three themes: unguided (free topic), value-guided, or controversy-guided. We filter out users who did not answer the preference survey or quit before completing all six conversations. As a result, we train on 1,000 users and evaluate zero-shot personalization performance on held-out users. This domain is particularly challenging due to the high heterogeneity and variance across users, as well as the lack of shared conversation topics, compared to the size of the training data. Despite the challenge, our work is among the first to report pluralistic reward modeling performance on the PRISM dataset.
5.2 Baselines
We evaluate PLUS against six existing RLHF baselines:
Bradley-Terry-Luce (BTL) model most widely used in RLHF preference learning [17]. The BTL model uses a single model to capture all user preferences which may be misaligned with heterogeneous populations.
Distributional Preference Learning (DPL)  Siththaranjan et al. [23] proposes to learn both the mean and the variance of the reward distribution, which captures overall variance across users’ preferences but does not model each user’s distinct preferences.
Variational Preference Learning (VPL)  Poddar et al. [18] learns a latent representation of each user with a vector of size 512 using a variational auto-encoder (VAE) reward model and shows state-of-the-art performance on pluralistic preference modeling.
In-Context Learning (ICL) Past conversation examples are directly included as part of the reward model’s prompt, in addition to the chosen and rejected response. This would in theory enable modeling each user’s distinct preferences, but may be less sample-efficient due to the lengthy context, and less robust to topic changes.
PLUS with a fixed summarizer (PLUS-untrained). This ablation only trains the reward model while fixing the summarizer.
Oracle. This is a reward model conditioned on the true user preferences. For example, in UltraFeedback, if a user values helpfulness and honesty over the other attributes, that user’s preference model is conditioned on the following string input: This user prefers helpfulness and honesty in the assistant’s responses.
@@ 6 Results
6.1 Advantages of PLUS on existing benchmarks
First, we compared PLUS against existing RLHF methods on existing personalized reward modeling benchmark tasks. Table 1 shows that PLUS achieves improved prediction accuracy across all three domains compared to vanilla RLHF with the largest improvement of 77% on Pets. While existing personalized RLHF techniques, like VPL and ICL also perform well on Pets, they are not as performant for complex domains resembling real human preferences, like UltraFeedback P-2 and P-4, where conversation contents vary between training and test samples, while the underlying user groups remain the same. VPL struggles to personalize showing limited improvement over BTL (-2.76–4.81% across model types). We hypothesize this is because training the reward model to distill all user context into a fixed-length vector sacrifices performance, as previously discussed by  Cideron et al. [4]. While in theory ICL can model each user’s distinct preferences, conditioning on all past conversation examples is sample-inefficient due to long, unsummarized contexts and is more sensitive to conversation topics than to the underlying user preferences. Specifically, ICL only shows 0.9–8.4% improvements across model types, compared to 10.4–17.7% achieved by PLUS, on the UltraFeedback dataset. We further investigate this limitation with our Pets-OOD experiments.
Table 1: Reward model accuracy on the held-out set. Mean and std from 3 seeds; otherwise, single seed.
@@ 6.1 Advantages of PLUS on existing benchmarks
BTL [2] 44.1 $\pm$ 2.8 58.1 $\pm$ 0.3 58.75 57.40 52.7 56.25 56.35
DPL [23] 46.2 $\pm$ 4.4 58.5 $\pm$ 0.7 58.15 57.55 56.2 57.00 56.60
VPL [18] 100 $\pm$ 0 58.4 $\pm$ 0.3 57.15 59.40 55.3 58.10 57.45
ICL 100 $\pm$ 0 59.6 $\pm$ 0.7 59.30 59.10 57.3 59.10 57.70
PLUS-untrained 96.2 $\pm$ 1.6 59.9 $\pm$ 0.2 59.65 59.35 56.4 58.75 57.90
PLUS (Ours) 99.8 $\pm$ 0.2 69.4 $\pm$ 1.1 69.85 66.90 61.8 62.45 62.80
Oracle 100 $\pm$ 0 69.0 $\pm$ 0.3 68.95 67.95 71.4 73.05 73.80
Even zero-shot summaries (by PLUS-untrained) can achieve improvements over BTL, but they still fall short of providing sufficient predictive signal for the reward model. This limitation is clearly observed in the Pets dataset: the untrained summarizer struggles to distinguish meaningful user contexts from spurious details or outputs vague descriptions, while the trained summarizer correctly identifies whether a user prefers cats or dogs, as shown in the box below. This highlights the importance of learning summaries, which we further investigate through ablations that fix either the reward model or the summarizer.
@@
6.2 Robustness of PLUS to new topics & users
As new users increasingly engage with LLM assistants, robustness to new users and conversation topics that are different from the training data becomes important for personalized systems. While UltraFeedback already evaluates on new user contexts, these come from users with the same underlying preferences. We took a step further by explicitly evaluating new users whose test-time preferences shift to entirely different content, using Pets (OOD). Concretely, the user context and preference labels $(s_{A},s_{B})$ used for training are about dogs or cats, but at test time, the summarizer receives a new user context about birds and rabbits, and the reward model is given two response options between birds and rabbits. As shown in Table 2, PLUS’s robust generalization to new user contexts indicates that it has learned a general algorithm for extracting user information into summaries, which translates well to new users with new types of preferences. In contrast, other approaches may be vulnerable to memorizing specific users’ preferences, rather than learning how to learn about what a new user is interested in.
Table 2: Reward model accuracy with and without user shift in Pets. The reward model is trained with Qwen2.5-0.5B-Instruct, and the summarizer is trained with Qwen2.5-3B-Instruct. Mean and std from 3 seeds. ICL’s performance drops from 100 to 72%, and VPL drops to below 50%, whereas PLUS is the only method to maintain accuracy above 90%.
Out-of-distribution user contexts
BTL 49.5 $\pm$ 2.6
DPL 49.0 $\pm$ 0.5
VPL 44.8 $\pm$ 4.0
ICL 72.5 $\pm$ 3.1
PLUS-untrained 85.8 $\pm$4.0
PLUS (Ours) 93.7 $\pm$4.0
[…]
6.3 Importance of co-adapting summarizer and reward
A key component of PLUS is the online co-adaptation of the summarizer and the reward model, which enables the summarizer to focus on dimensions of user variability that are most relevant to the reward model. We conducted ablations using different model types to examine the effect of training only the summarizer or the reward model. Table 3 shows that despite using a smaller reward model (0.5B), co-adaptation achieves higher performance than larger reward models (3B) that are either trained with fixed summaries or only used to train the summarizer. Surprisingly, using a state-of-the-art summarizer (GPT-4o) with a trained 3B reward model still underperforms compared to co-training substantially smaller models (0.5B reward model and a 3B summarizer) in both UltraFeedback domains. This supports our intuition about the importance of RL fine-tuning for learning summaries: good summaries are defined by the reward model’s performance rather than any fixed criteria. Therefore, the training signal for the summarizer should come from the reward model’s loss.
@@ 6.3 Importance of co-adapting summarizer and reward
6.4 Flexibility to incorporate unstructured text data beyond preference labels
@@
Table 4: Reward model accuracy with unstructured inputs. We compared the performance of PLUS and ICL when user contexts include unstructured data other than preference labels, as most LLM-user interactions do not gather explicit preference feedback. (i) Preference data provides users’ past preferences in the form of chosen and rejected response examples. (ii) User-guideline + preference data additionally includes textual instructions specifying which aspects of an LLM assistant the summarizer should focus on. (iii) Conversation data does not include binary preference labels; instead, it only presents the chosen responses as “examples of past conversation” and expects the reward model or summarizer to infer the user’s underlying preferences only from the positive examples. This last setup more closely models real user interactions with LLM assistants, since users are typically given only one response rather than choosing between two options; however, their preferences can still be inferred from the users’ initial prompts, follow-up responses, and questions. We trained a Qwen2.5-0.5B-Instruct reward model for both ICL and PLUS and a Qwen2.5-3B-Instruct summarizer for PLUS. We did not include VPL, as it only supports pairwise preferences as contexts.
@@
6.5 Evaluating PLUS in the wild with the PRISM user data
Despite the growing interest in pluralistic alignment, results demonstrating effectiveness on real-world datasets have been limited due to both the challenging nature of real user data and the limited sample size. Having established the advantages of PLUS on existing benchmarks for pluralistic alignment, we next applied it to the real-user dataset PRISM [12].
Reward modeling with PRISM. We first implemented PLUS with a Qwen2.5-3B-Instruct summarizer and a Qwen2.5-0.5B-Instruct reward model using 1,000 users from the PRISM dataset. We also implemented six reward-model baselines with Qwen2.5-0.5B-Instruct. Then, we evaluated reward-model accuracy on the preference labels of held-out conversations. This evaluation protocol assumes that each user’s underlying preferences remain consistent across different conversation topics, so once the reward model learns a user’s preferences from five conversations, it should be able to predict the preferred response in a new, sixth conversation. Due to the challenging nature of this dataset, which attempts to capture a high degree of variance with a small sample size, we find that all methods achieve somewhat low reward prediction accuracy, and report the results in Table 6 in the appendix. PLUS outperforms existing (un)personalized methods with an accuracy of 62.9%, whereas BTL achieves 59.8%, VPL 60.4%, and ICL 60.1%. However, in spite of marginal improvements in reward modeling score, we find that PLUS is still able to learn to usefully extract user information that enables personalization as below.
6.6 Zero-shot personalization on PRISM with GPT-4 models using PLUS summaries
So far, we have focused on the reward modeling performance of PLUS compared to existing RLHF techniques. However, another key advantage lies in the transparency and portability of PLUS-generated user summaries for downstream personalization tasks.
Personalized LLM-as-judges. We combine PLUS with an LLM-as-judge setup [33] to enable zero-shot personalized reward modeling. Specifically, we evaluate GPT-4’s prediction of a user’s preferred response between two candidates, both with and without the PLUS-generated summary. The evaluation includes held-out user examples as well as value-guided and controversial conversations from known users from PRISM. As shown in Table 7, across three task types, PLUS-guided LLM-as-judges improve accuracy for both GPT-4o (by 19.5%) and GPT-4.1 (by 2.9%). These results demonstrate that PLUS-generated summaries are effective not only for personalizing learned reward models but also for personalizing zero-shot proprietary models.
Personalized response generation.
We prompted GPT-4o and 4.1 to generate responses with and without the PLUS-generated user summary. Specifically, we sampled two new prompts for each of the 308 held-out users in PRISM and compared the win rates of personalized versus default GPT responses evaluated by an oracle reward model, which is trained to condition on users’ self-stated preferences (in PRISM, users rate different attributes of an LLM assistant in surveys, which we use as ground truth). This experiment assesses whether PLUS-generated summaries are useful for enabling proprietary models to personalize to new users without additional finetuning. In Fig. 3, the win rates of PLUS-personalized responses over default GPT responses clearly demonstrate the effectiveness of PLUS (e.g., 72% vs. 28% for GPT-4o, and 69% vs. 31% for GPT-4.1). We included examples below to show the effect of PLUS-summary-conditioned personalization using GPT-4.
7 Discussion
We propose PLUS, a novel RLHF technique of learning text-based summaries to capture user variability and co-training a reward model conditioned on the obtained summaries. Summaries provide a natural format that leverages the language modeling capabilities of LLMs and show strong empirical performance that extends beyond benchmark tasks. We conducted extensive experiments and compared our approach against six RLHF baselines, demonstrating the following five benefits.
  • Strong empirical advantages. PLUS outperforms standard BTL by 11-77% on pluralistic datasets of Pets and UltraFeedback, and achieves consistently strong performance compared to existing pluralistic RLHF techniques.
  • Robustness. PLUS is the only method that maintains 90% accuracy with test-time preference content shifts, suggesting that it has learned a general algorithm for extracting user information that translates to new users with new types of preferences.
  • Benefit of co-adaptation. Co-adapting a smaller summarizer (3B parameters) and reward model (0.5B) outperforms a reward model trained on proprietary model summaries by 10%. We carefully compared PLUS to alternative implementations of training either the summarizer or the reward model while fixing the other. Our experiments show that automatic summaries affect the learnability of the reward model, and only training the summarizer achieves poor reward modeling performance across domains and model types.
  • Flexibility to incorporate unstructured user contexts. Unlike prior personalized RLHF techniques [18, 14] , PLUS can leverage user-guided instructions and unlabeled conversation data, achieving improvements of 9-16% compared to ICL. This is especially important for scaling up to real user systems where explicit preference labels may not be available, and therefore the user contexts must be inferred from other types of text data.
  • Summary-conditioned personalization of proprietary models, achieving a win rate of 72% against the unpersonalized GPT-4o responses. Furthermore, generated summaries are human-readable, enhancing transparency and improving users’ trust in LLM systems.
Appendix
Appendix A Illustrative example of why compression by $z$ matters.
Naively summarizing past conversations or using the entire context $c$ may be insufficient. Summarizing past conversations with an untrained summarizer can lead to summaries that focus on past content rather than the general style of conversations the user prefers to have. Therefore, the obtained summaries may not provide predictive signal for the reward model.
Figure 4: An intuitive example of why compression by $z$ matters for the reward model.
As an illustrative example, in Fig. 4, suppose the reward is +1 for the area above $y=x$ and -1 otherwise. The objective is to represent the context vector $c$ as a scalar. A lossless compression outputs $z$ as the angle between $c$ and the diagonal, while an uninformed summarizer may instead output $z^{\prime}=\theta_{1}$. This affects the expressivity of the learned reward model, since this lossy compression can never recover the correct reward.
Alternatively, one could include all prior preference labels from the user as in-context learning (ICL) examples, and train the reward model conditioned on such examples (i.e., the full list of examples acting as the conditioning variable $z$). However, the high-dimensional input space of ICL, and subsequently a more complex hypothesis class of reward models, may make reward learning more difficult in terms of sample complexity. This shows up in our empirical results, where ICL underperforms when trying to generalize to new conversation topics. While in theory ICL can enable personalization, it may be sample-inefficient and computationally expensive to train and use for inference due to the increased context.
Appendix B PLUS-generated summary examples

### v3

@@ 5 Experiments
Domains. We evaluate on three benchmark tasks from  Poddar et al. (2024), along with a challenging variant that tests out-of-distribution user contexts, as well as real user datasets. Pets. In this dataset, 50% of users prefer to discuss dogs and 50% prefer cats, so a single reward model will fail to accurately capture user preferences. To test the robustness of different methods with new conversation topics and users, we create a new Pets OOD (Out-Of-Distribution) dataset, which consists of users who prefer talking about rabbits or birds, as opposed to cats and dogs. Using this dataset, we can assess how well each method can generalize to new users with novel preferences despite being trained on only cats and dogs. Poddar et al. (2024) construct two domains from the scaled-up preference dataset UltraFeedback (Cui et al., 2024). In UltraFeedback-P2, users prefer either helpfulness or honesty, and in UltraFeedback-P4, users choose a subset of four possible attributes of an LLM assistant: helpfulness, honesty, truthfulness, and instruction-following. In both domains, personalized reward models need to infer different user types based on past conversation data. For example, some users may prioritize helpfulness and honesty, while others may value instruction-following above all other attributes. Following  Poddar et al. (2024), we also focus on a controversial subset of the dataset by removing samples where all users agree on the chosen response. Evaluation with real user data (PRISM). This real-world user dataset consists of open-ended conversations from 1,500 participants across 75 countries and 20 LLMs (Kirk et al., 2024). Each user engages in up to six conversations on topics of their choice, loosely guided by one of three themes: unguided (free topic), value-guided, or controversy-guided. We filter out users who did not answer the preference survey or quit before completing all six conversations. As a result, we train on 1,000 users and evaluate personalization performance on held-out conversation topics and/or users. This domain is particularly challenging due to the high heterogeneity and variance across users, as well as the lack of shared conversation topics, compared to the size of the training data. Despite the challenge, our work is among the first to report pluralistic reward modeling performance on the PRISM dataset.
Baselines. We evaluate PLUS against six existing RLHF baselines: Bradley-Terry-Luce (BTL) model most widely used in RLHF preference learning Ouyang et al. (2022). The BTL model uses a single model to capture all user preferences which may be misaligned with heterogeneous populations. Distributional Preference Learning (DPL)  Siththaranjan et al. (2024) proposes to learn both the mean and the variance of the reward distribution, which captures overall variance across users’ preferences but does not model each user’s distinct preferences. Variational Preference Learning (VPL)  Poddar et al. (2024) learns a latent representation of each user with a vector of size 512 using a variational auto-encoder (VAE) reward model and shows state-of-the-art performance on pluralistic preference modeling. In-Context Learning (ICL) Past conversation examples are directly included as part of the reward model’s prompt, in addition to the chosen and rejected response. This would in theory enable modeling each user’s distinct preferences, but may be less sample-efficient due to the lengthy context, and less robust to topic changes. PLUS with a fixed summarizer (PLUS-untrained). This ablation only trains the reward model while fixing the summarizer. Oracle. This is a reward model conditioned on the true user preferences. For example, in UltraFeedback, if a user values helpfulness and honesty over the other attributes, that user’s preference model is conditioned on the following string input: This user prefers helpfulness and honesty in the assistant’s responses.
@@ 6 Results
Advantages of PLUS on existing benchmarks.
First, we compared PLUS against existing RLHF methods on existing personalized reward modeling benchmark tasks. Table 1 shows that PLUS achieves improved prediction accuracy across all three domains compared to vanilla RLHF with the largest improvement of 77% on Pets. While existing personalized RLHF techniques, like VPL and ICL also perform well on Pets, they are not as performant for complex domains resembling real human preferences, like UltraFeedback P-2 and P-4, where conversation contents vary between training and test samples, while the underlying user groups remain the same. VPL struggles to personalize showing limited improvement over BTL (-2.76–4.81% across model types). We hypothesize this is because training the reward model to distill all user context into a fixed-length vector sacrifices performance, as previously discussed by  Cideron et al. (2022). While in theory ICL can model each user’s distinct preferences, conditioning on all past conversation examples is sample-inefficient due to long, unsummarized contexts and is more sensitive to conversation topics than to the underlying user preferences. Specifically, ICL only shows 0.9–8.4% improvements across model types, compared to 10.4–17.7% achieved by PLUS, on the UltraFeedback dataset. Even for larger reward models (Llama-8B-Instruct and Qwen-7B-Instruct), ICL still perform less accurately than 3B PLUS for Ultrafeedback domains (Table 6). We further investigate this limitation with our Pets-OOD experiments.
Table 1: Reward model accuracy on the held-out set. We report the mean and two-sided 95% t-intervals over 3 seeds; otherwise, a single seed.
@@
BTL (Bradley and Terry, 1952) 44.15 $\pm 7.02$ 58.12 $\pm$ 0.87 58.75 57.40 53.53 $\pm$ 3.27 56.25 56.35
DPL Siththaranjan et al. (2024) 46.20 $\pm$ 10.95 58.53 $\pm$ 1.74 58.15 57.55 55.22 $\pm$ 3.71 57.00 56.60
VPL Poddar et al. (2024) 100 $\pm$ 0 58.37 $\pm$ 0.87 57.15 59.40 54.2 $\pm$ 5.84 58.10 57.45
ICL 100 $\pm$ 0 59.62 $\pm$ 1.79 59.30 59.10 57.22 $\pm$ 0.95 59.10 57.70
PLUS-untrained 96.20 $\pm$ 4.13 59.90 $\pm$ 0.50 59.65 59.35 56.89 $\pm$ 1.11 58.75 57.90
PLUS (Ours) 99.80 $\pm$ 0.86 69.40 $\pm$ 2.81 69.85 66.90 62.10 $\pm$ 0.69 62.45 62.80
Oracle 100 $\pm$ 0 69.10 $\pm$ 0.72 68.95 67.95 70.60 $\pm$ 1.93 73.05 73.80
Even summaries by PLUS-untrained can achieve improvements over BTL, but they still fall short of providing sufficient predictive signal for the reward model. This limitation is clearly observed in the Pets dataset: the untrained summarizer struggles to distinguish meaningful user contexts from spurious details or outputs vague descriptions, while the trained summarizer correctly identifies whether a user prefers cats or dogs, as shown in the box below. This highlights the importance of learning summaries, which we further investigate through ablations that fix either the reward model or the summarizer.
Robustness of PLUS to new topics & users.
As new users increasingly engage with LLM assistants, robustness to new users and conversation topics that are different from the training data becomes important for personalized systems. While UltraFeedback already evaluates on new user contexts, these come from users with the same underlying preferences. We took a step further by explicitly evaluating new users whose test-time preferences shift to entirely different content, using Pets (OOD).
Concretely, the user context and preference labels $(s_{A},s_{B})$ used for training are about dogs or cats, but at test time, the summarizer receives a new user context about birds and rabbits, and the reward model is given two response options between birds and rabbits. As shown in Table 2, PLUS’s robust generalization to new user contexts indicates that it has learned a general algorithm for extracting user information into summaries, which translates well to new users with new types of preferences. In contrast, other approaches may be vulnerable to memorizing specific users’ preferences, rather than learning how to learn about what a new user is interested in.
Table 2: Reward model accuracy with and without user shift in Pets. We report the mean and the two-sided 95% t-intervals for 3 seeds. The reward model is trained with Qwen2.5-0.5B-Instruct, and the summarizer is trained with Qwen2.5-3B-Instruct. Mean and std from 3 seeds. ICL’s performance drops from 100 to 72%, and VPL drops to below 50%, whereas PLUS is the only method to maintain accuracy above 90%.
Out-of-distribution user contexts
BTL 49.5 $\pm$ 6.45
DPL 49.0 $\pm$ 1.24
VPL 44.83 $\pm$ 10.04
ICL 72.50 $\pm$ 7.76
PLUS-untrained 85.83 $\pm$ 10.04
PLUS (Ours) 93.67 $\pm$ 9.96
[…]
@@
Importance of co-adapting summarizer and reward.
A key component of PLUS is the online co-adaptation of the summarizer and the reward model, which enables the summarizer to focus on dimensions of user variability that are most relevant to the reward model. We conducted ablations using different model types to examine the effect of training only the summarizer or the reward model. Table 3 shows that despite using a smaller reward model (0.5B), co-adaptation achieves higher performance than larger reward models (3B) that are either trained with fixed summaries or only used to train the summarizer. Surprisingly, using a strong, proprietary summarizer (GPT-4o) with a trained 3B reward model still underperforms compared to co-training substantially smaller models (0.5B reward model and a 3B summarizer) in both UltraFeedback domains. This supports our intuition about the importance of RL fine-tuning for learning summaries: good summaries are defined by the reward model’s performance rather than any fixed criteria. Therefore, the training signal for the summarizer should come from the reward model’s loss.
@@
Incorporating flexible textual inputs beyond preference data.
@@
Table 4: Reward model accuracy with unstructured inputs. The reward model is trained with Qwen2.5-0.5B-Instruct, and the summarizer is trained with Qwen2.5-3B-Instruct on the UltraFeedback datasets.
UF-P2 Accuracy
Non-preference User guide
ICL 58.55 59.65
PLUS 68.30 70.25
UF-P4 Accuracy
Non-preference User guide
ICL 56.96 56.60
PLUS 62.45 62.70
Figure 3: PPO training return curves for Pets, Ultrafeedback P2, and PRISM using Qwen3B-Instruct summarizer and Qwen0.5B reward model with a rollout batch size 256. The returns obtained by the summarizer are the negative prediction (log-likelihood) loss of the reward model.
Evaluating PLUS in the wild with PRISM user data. Despite the growing interest in pluralistic alignment, results demonstrating effectiveness on real-world datasets have been limited due to both the challenging nature of real user data and the limited sample size. Having established the advantages of PLUS on existing benchmarks for pluralistic alignment, we next applied it to the real-user dataset PRISM (Kirk et al., 2024). Reward modeling with PRISM. We first implemented PLUS with a Qwen2.5-3B-Instruct summarizer and a Qwen2.5-0.5B-Instruct reward model using 1,000 users from the PRISM dataset. We also implemented six reward-model baselines with Qwen2.5-0.5B-Instruct. Then, we evaluated reward-model accuracy on the preference labels of held-out conversations. This evaluation protocol assumes that each user’s underlying preferences remain consistent across different conversation topics, so once the reward model learns a user’s preferences from five conversations, it should be able to predict the preferred response in a new, sixth conversation. Due to the challenging nature of this dataset, which attempts to capture a high degree of variance with a small sample size, we find that all methods achieve somewhat low reward prediction accuracy, and report the results in Table 8 in the appendix (in particular, when used to predict out-of-distribution users, the pluralistic reward models, e.g., VPL, ICL and PLUS, perform poorly  11). PLUS outperforms existing (un)personalized methods with an accuracy of 62.9%, whereas BTL achieves 59.8%, VPL 60.4%, and ICL 60.1% for the in-distribution users. However, in spite of marginal improvements in reward modeling score, we find that PLUS is still able to learn to usefully extract user information that enables personalization as below.
Personalization on PRISM with GPT-4 models using PLUS summaries without fine-tuning. So far, we have focused on the reward modeling performance of PLUS compared to existing RLHF techniques. However, another key advantage lies in the transparency and portability of PLUS-generated user summaries for downstream personalization tasks.
Personalized LLM-as-judges. We combine PLUS with an LLM-as-judge setup (Zheng et al., 2023) to enable personalized reward modeling without further fine-tuning. Specifically, we evaluate GPT-4’s prediction of a user’s preferred response between two candidates, both with and without the PLUS-generated summary. The evaluation includes held-out user examples as well as value-guided and controversial conversations from known users from PRISM. As shown in Table 9, across three task types, PLUS-guided LLM-as-judges improve accuracy for both GPT-4o (by 19.5%) and GPT-4.1 (by 2.9%). These results demonstrate that PLUS-generated summaries are effective not only for personalizing learned reward models but also for personalizing proprietary models without further fine-tuning.
Personalized response generation. We prompted GPT-4o and 4.1 to generate responses with and without the PLUS-generated user summary. Specifically, we sampled two new prompts for each of the 308 held-out users in PRISM and compared the win rates of personalized versus default GPT responses evaluated by an oracle reward model, which is trained to condition on users’ self-stated preferences (in PRISM, users rate different attributes of an LLM assistant in surveys, which we use as ground truth). This experiment assesses whether PLUS-generated summaries are useful for enabling proprietary models to personalize to new users without additional finetuning. In Fig. 4, the win rates of PLUS-personalized responses over default GPT responses clearly demonstrate the effectiveness of PLUS (e.g., 72% vs. 28% for GPT-4o, and 69% vs. 31% for GPT-4.1). We included examples in Appendix F to show the effect of PLUS-summary-conditioned personalization using GPT-4.
7 Discussion & Limitations
We propose PLUS, a novel RLHF technique of learning text-based summaries to capture user variability and co-training a reward model conditioned on the obtained summaries. Summaries provide a natural format that leverages the language modeling capabilities of LLMs and show strong empirical performance that extends beyond benchmark tasks. We conducted extensive experiments and compared our approach against six RLHF baselines, demonstrating the following five benefits. Strong empirical advantages. PLUS outperforms standard BTL by 11-77% on pluralistic datasets of Pets and UltraFeedback. Robustness. PLUS is the only method that maintains 90% accuracy with test-time preference content shifts. Benefit of co-adaptation. Co-adapting a smaller summarizer (3B parameters) and reward model (0.5B) outperforms a reward model trained on proprietary model summaries by 10%. Flexibility to incorporate unstructured user contexts. Unlike prior personalized RLHF techniques (Poddar et al., 2024; Li et al., 2024) , PLUS can leverage user-guided instructions and unlabeled conversation data, achieving improvements of 9-16% compared to ICL. Summary-conditioned personalization of proprietary models, achieving a win rate of 72% against the unpersonalized GPT-4o responses. Furthermore, generated summaries are human-readable, enhancing transparency and improving users’ trust in LLM systems. Limitations. PPO in a MARL setup, which enables co-adaptation of the summarizer and the reward model, may be sensitive to training hyperparameters, such as learning rates. Some potential approaches to mitigate training issues (e.g., instability, under-convergence) may include training the summarizer and reward model for more than one epoch, as observed by  Poddar et al. (2024) (for their latent encoder and the reward model); or specifically, for our MARL setup, freezing the weights of the summarizer after one epoch and continuing to train only the reward model conditioned on the generated summaries, or warm-starting the reward model by first training it on the outputs of the untrained summarizer. These approaches may be helpful while taking care to avoid overfitting to the training data.
Ethics Statement. We propose PLUS as a step toward pluralistic preference alignment that is both interpretable and accessible even for non-technical users, as the user representation is in human-readable text form. We hope to broaden the benefits of personalized and value-aligned LLM assistants to a wider audience, while being cognizant of ethical considerations, like privacy and user trust in AI systems. Privacy considerations. We believe that any personalized reward modeling methods that require maintaining information for each user or group of users may run into the issue of using a user’s private data to distinguish between different users. Therefore, it would be crucial to consider removing any sensitive information prior to training a user-conditioned reward model (ICL, VPL, PLUS) or a PLUS summarizer for actual deployment on users. Given the importance of transparency and maintaining user trust in personalized AI systems, we believe that the ability to present user summaries in natural language is a comparative advantage of PLUS that other personalized reward models do not afford. The user-specific summaries can be shown to the user to reveal which parts of the previous conversation are being prioritized by the reward model, and it would be interesting to consider allowing users to edit or remove specific information from these summaries.
Reproducibility. Our code is available at this link. The co-adaptation pipeline in Alg. 1 was built on top of PPO training framework from OpenRLHF (Hu et al., 2024). We also used the reward model training code from OpenRLHF. For other baselines, namely VPL, DPL, and BTL (for Pets), we used the implementations by  Poddar et al. (2024). Our experiments were conducted with 4 $\times$ H200 GPUs.
Appendix
Appendix A Inferring user contexts with flexible textual inputs
@@
Table 5: Reward model accuracy with unstructured inputs. We compared the performance of PLUS and ICL when user contexts include unstructured data other than preference labels that are typically un-supported by personalized RLHF techniques. Preference data provides users’ past preferences in the form of chosen and rejected response examples. User-guideline + preference data additionally includes textual instructions specifying which aspects of an LLM assistant to focus on. Conversation data does not include binary preference labels; instead, it only presents the chosen responses as “examples of past conversation” and expects the reward model or summarizer to infer the user’s underlying preferences only from the positive examples. This last setup more closely models real user interactions with LLM assistants, since users are typically given only one response rather than choosing between two; however, their preferences can still be inferred from their questions, initial prompts, and follow-up responses, despite not having observed explicit user preferences. We trained a Qwen2.5-0.5B-Instruct reward model for both ICL and PLUS and a Qwen2.5-3B-Instruct summarizer for PLUS. Interestingly, PLUS benefits from additional user instructions specifying the relevant dimensions of an LLM assistant’s attributes, whereas ICL’s performance is hurt by longer user prompts. We did not compare with VPL, as it does not support unstructured input formats.
@@
Appendix B Performance of ICL with larger reward models
Table 6: ICL reward model accuracy with larger base model sizes. We included 7B & 8B reward models for UF-P2 and UF-P4 and observed that the performance achieved by these larger ICL models is lower than the performance obtained by PLUS with a 3B reward model and a 3B summarizer (66.9 on UF-P2 and 62.8 on UF-P4 with Llama-3B).
Preference Prediction Accuracy
Base model Ultrafeedback-P2 Ultrafeedback-P4
Llama8B-Instruct 61.4 58.6
Qwen7B-Instruct 61.8 58.8
Appendix C PLUS-generated summary examples